During a coffee-room debate among several young M.B.A.s who had recently graduated, one of the young executives flatly stated, “The most this company can lose on its Brazilian division is the amount it has invested (its fixed costs).” Not everyone agreed with this statement. In what sense is this statement correct? Under what circumstances could it be false? Explain


1
Expert's answer
2021-01-15T18:21:01-0500

The statement is partially correct. This is because the firm loses fixed cost only when it is in the stage of shutting down.

- The condition for firms to shut down is when the total revenue becomes less than the total variable cost. Then the firm loses only fixed cost.

- However if the firm is getting lower total revenue than total cost, but it is not less than the total variable cost, then firms should operate to stop making larger losses. For example, if total revenue is $40, the total variable cost is $10 and the total fixed cost is $100, then profit is equal to:

"Profit = TR -TVC -TFC \\

= 40 -10 -100 \\

= -70"

- If the firm stops its unit in this stage, TR and TVC would be equal to zero as Q = 0. The losses incurred would be $100 of fixed cost. In this case, the possible loss could have been -70.